Work out the specific heat capacity of a substance if it takes 5200J of energy to heat 2kg of it from 42°C to 47°C.
1
Expert's answer
2020-03-03T10:03:50-0500

"c=\frac {\Delta \Epsilon}{m\Delta\theta}"

Where c is the specific heat capacity(J/Kg °C)

ΔE is change in the thermal energy(J)

m is  mass (kg)

 Δθ change in temperature(°C)

"=\frac {5200J} {2kg *(47-47)C}"

=520J/kg°C